OpenStudy (anonymous):

If M,A,T,H,S denote +ve Real nos. such that MA=12, TH=30, AH=24, AT=20, HS=42. Find MATHS

OpenStudy (anonymous):

2520

OpenStudy (anonymous):

M = 3, A = 4, T = 5, H = 6, S = 7

OpenStudy (anonymous):

H = 30/T T = 20/A therefore: H = (3/2)A therefore: (3/2)A^2 = 24 therefore: A = 4 The rest is really easy.

OpenStudy (anonymous):

Then the answer is 1890 rather than 2520 @MrMoose.

OpenStudy (anonymous):

No, it is 2520

OpenStudy (anonymous):

Hey, which answer is corect.... please solve it corectly and reply..... @MrMoose @Zekarias

OpenStudy (mayankdevnani):

hey! you just really do multiply

OpenStudy (mayankdevnani):

23*30*24*20*42=13910400

OpenStudy (anonymous):

how did u get that 23... @mayankdevnani

OpenStudy (mayankdevnani):

sry.. 24

OpenStudy (anonymous):

no... u r doing a bit of mistake

OpenStudy (mayankdevnani):

12*30*24*20*42=7257600

OpenStudy (mayankdevnani):

ok... sry

OpenStudy (anonymous):

it is d value of MATHAHATHS... but i need value of MATHS.... @mayankdevnani